OpenStudy (anonymous):

7/10c - 10 = 2/5 Find c.

OpenStudy (driftracer305):

k so 2/5 + 10 = 7/10c

OpenStudy (driftracer305):

then 52/5 = 7/10c

OpenStudy (driftracer305):

so, 7*5 = 520*c

OpenStudy (nick88888888):

@driftracer305 Oh i solved it differently

OpenStudy (driftracer305):

yea sure @Nick88888888 ....... there are multiple ways u can solve this.....

OpenStudy (anonymous):

nick can you solve the problem step by step?

OpenStudy (driftracer305):

@xLogic .. he is gone buddy

OpenStudy (nick88888888):

No im here

OpenStudy (driftracer305):

k solve it for him.... i dont think he gets my simple way....

OpenStudy (nick88888888):

7/10 c - 10 = 2/5 7/2*5 c - 10 = 2/5 7c/ 2*5 - 10 = 2/5 c = 104/7 thats how i did it but.... derp